### v1

@@ Abstract
Online construction of open-ended language scenes is crucial for robotic applications, where open-vocabulary interactive scene understanding is required. Recently, neural implicit representation has provided a promising direction for online interactive mapping. However, implementing open-vocabulary scene understanding capability into online neural implicit mapping still faces three challenges: lack of local scene updating ability, blurry spatial hierarchical semantic segmentation and difficulty in maintaining multi-view consistency. To this end, we proposed O2V-mapping, which utilizes voxel-based language and geometric features to create an open-vocabulary field, thus allowing for local updates during online training process. Additionally, we leverage a foundational model for image segmentation to extract language features on object-level entities, achieving clear segmentation boundaries and hierarchical semantic features. For the purpose of preserving consistency in 3D object properties across different viewpoints, we propose a spatial adaptive voxel adjustment mechanism and a multi-view weight selection method. Extensive experiments on open-vocabulary object localization and semantic segmentation demonstrate that O2V-mapping achieves online construction of language scenes while enhancing accuracy, outperforming the previous SOTA method.
@@ 1 Introduction
Nonetheless, aforementioned methods merely utilize heavy MLPs to represent the scene, leading to lengthy training time and they can only operates off-line on pre-collected data. Moreover, they introduce additional decoder branches to output semantic labels which adds extra burden to the framework. To adapt to complex and rapidly changing scenes, an online open-vocabulary mapping framework is significantly needed.
In terms of open-set semantic scene understanding, the common problem of lies in the semantic information varying depending on different scales. For example, the object can be interpreted as the door and the door handle as its focus changes. Additionally, semantic vagueness is widely observed, where clear semantic boundaries cannot be obtained at the edges of objects. The key reason is that most existing works focus on pixel-level semantics, lacking object-level semantic understanding. Furthermore, during the process of online mapping, observing the same object from different perspectives may result in the assignment of different semantics to that object. The issues mentioned above contribute to spatiotemporal ambiguity in the process of semantic mapping.
In this paper, we present O2V-Mapping, a novel method for online constructing open-set semantic scenes efficiently and accurately. Our framework achieves online open-vocabulary mapping by grounding language embeddings from text-image large models like CLIP into voxel-based neural implicit representation. Moreover, by reconstructing open-set semantic scenes at the object level and utilize segmentation priors of foundation models, we eliminate the semantic vagueness, get distinct semantic boundaries and resolve the multi-view inconsistency. Finally, we propose a LLM-centric agent architecture, instantiating our open-set semantic scene approach as the interactive memory, enabling full-scene grounding through query and render mechanisms to implement grounded tree search and online memory refinement. The main contributions of this work are summarized as follows.
@@ 2 Related Work
Open-vocabulary 3D scene understanding. Recent advances in the field of open-vocabulary 3D scene understanding has been marked by the confluence of 2D visual language models and 3D point cloud processing techniques. The advent of Contrastive Vision-Language Pre-training (CLIP)[18] has notably enhanced the ability to comprehend scenes in a 2D open-vocabulary context without the need for semantic labels [6, 9, 13, 14, 26, 27, 28, 32, 33, 35, 36]. Several researchers have effectively aligned the textual and visual features derived from CLIP with 3D point cloud features, facilitating the expansion of open-vocabulary scene understanding to 3D scenes. Tschernezki et al. [24] introduced the DINO model [12] for extracting 2D features, which are then subsequently utilized for semantic feature learning within a trainable semantic field. Furthermore, Segment Anything Model (SAM) [5] has demonstrated remarkable zero-shot capabilities across a multitude of downstream 2D tasks. Extending this innovation, SAM3D [30] integrates the 2D masks generated by SAM with point clouds through projection, thereby imbuing the 3D point cloud space with semantic features.
@@ 3.1 Online Construction of Open-Vocabulary Field
Current methods about constructing dense open-set semantic scenes such as LERF [4] mainly rely on resource-intensive neural networks to embed CLIP features. This restricts their ability to be updated online. To build an online open-vocabulary mapping pipeline, we employ hierarchical scene representations of various scales and types to store geometric and semantic information about scene objects. Specifically, the scene is represented by three types of voxel grids: RGB, geometric, and semantic. These three distinct features are associated to each cell of the feature grid and updated in on online manner. We utilize Multilayer Perceptrons (MLPs) to decode occupancy, color and semantic information from these features and perform volume rendering. Specifically, for a spatial point $p$, we extract two fused latent vectors: the depth feature ${\phi}_{d}(p)$ and the color feature ${\phi}_{c}(p)$. We use trilinear interpolation between the eight neighboring voxels to extract latent vectors. These vectors are concatenated with the position-encoded point $p$ and fed into two separate MLPs, $f_{d}$ and $f_{c}$, to decode occupancy probability and color values, respectively.
@@ 4.1 Qualitative Results
O2V-mapping not only allows for the online construction of open-set vocabulary scenes but also significantly improves object-level scene understanding, semantic segmentation accuracy, and precision compared to LERF. For larger objects like "couch" that only appear in small portions in each image, O2V-mapping can still leverage the spatial consistency of implicit scene representations to understand the object as a whole, and demonstrate complete and clear object boundaries. For small yet complex objects like "potted plant", O2V-mapping can also describe precise object boundaries. Additionally, multiple objects of the same type, such as "chair" and "trash can" in the scene can also be assigned correct open-set semantic information by O2V-mapping. Even with slight differences in material, color, and shape among objects of the same type, they can still be clearly distinguished from surrounding objects of different types. For objects like "table" and "door" with uniform textures and flat surfaces in the scene, O2V-mapping exhibits excellent performance.
@@ 4.2 Query Robustness
LERF is highly sensitive to query keywords, making it prone to false positives when the query keywords are ambiguous or exhibit a "bag-of-words" behavior and struggles to capture spatial relationships between objects. A common scenario is that when the queried object is not within the current viewpoint, LERF tends to exhibit a confused distribution of relevance scores (5). O2V-mapping exhibits robustness to query keywords. As shown in 5, in scenes where "door" is not present within the current viewpoint, there is no confused distribution of relevance scores.

### v2

@@ Abstract
Online construction of open-ended language scenes is crucial for robotic applications, where open-vocabulary interactive scene understanding is required. Recently, neural implicit representation has provided a promising direction for online interactive mapping. However, implementing open-vocabulary scene understanding capability into online neural implicit mapping still faces three challenges: lack of local scene updating ability, blurry spatial hierarchical semantic segmentation and difficulty in maintaining multi-view consistency. To this end, we proposed O2V-mapping, which utilizes voxel-based language and geometric features to create an open-vocabulary field, thus allowing for local updates during online training process. Additionally, we leverage a foundational model for image segmentation to extract language features on object-level entities, achieving clear segmentation boundaries and hierarchical semantic features. For the purpose of preserving consistency in 3D object properties across different viewpoints, we propose a spatial adaptive voxel adjustment mechanism and a multi-view weight selection method. Extensive experiments on open-vocabulary object localization and semantic segmentation demonstrate that O2V-mapping achieves online construction of language scenes while enhancing accuracy, outperforming the previous SOTA method.We have now open-sourced our code at https://github.com/Fudan-MAGIC-Lab/O2Vmapping.git.
@@ 1 Introduction
Nonetheless, aforementioned methods merely utilize heavy MLPs to represent the scene, leading to lengthy training time and they can only operate off-line on pre-collected data. Moreover, they introduce additional decoder branches to output semantic labels which adds extra burden to the framework. To adapt to complex and rapidly changing scenes, an online open-vocabulary mapping framework is significantly needed.
In terms of open-set semantic scene understanding, the common problem lies in the semantic information varying depending on different scales. For example, the object can be interpreted as the door and the door handle as our focus changes. Additionally, semantic vagueness is widely observed, where clear semantic boundaries cannot be obtained at the edges of objects. The key reason is that most existing works focus on pixel-level semantics, lacking object-level semantic understanding. Furthermore, during the process of online mapping, observing the same object from different perspectives may result in the assignment of different semantics to that object. The issues mentioned above contribute to spatiotemporal ambiguity in the process of semantic mapping.
In this paper, we present O2V-Mapping, a novel method for online constructing open-set semantic scenes efficiently and accurately. Our framework achieves online open-vocabulary mapping by grounding language embeddings from text-image large models like CLIP into voxel-based neural implicit representation. Moreover, by reconstructing open-set semantic scenes at the object level and utilizing segmentation priors of foundation models, we eliminate the semantic vagueness, get distinct semantic boundaries and resolve the multi-view inconsistency. Finally, we propose a LLM-centric agent architecture, instantiating our open-set semantic scene approach as the interactive memory, enabling full-scene grounding through query and render mechanisms to implement grounded tree search and online memory refinement. The main contributions of this work are summarized as follows.
@@ 2 Related Work
Open-vocabulary 3D scene understanding. Recent advances in the field of open-vocabulary 3D scene understanding has been marked by the confluence of 2D visual language models and 3D point cloud processing techniques. The advent of Contrastive Vision-Language Pre-training (CLIP)[19] has notably enhanced the ability to comprehend scenes in a 2D open-vocabulary context without the need for semantic labels [7, 10, 14, 15, 27, 28, 29, 33, 34, 36, 37]. Several researchers[3] have effectively aligned the textual and visual features derived from CLIP with 3D point cloud features, facilitating the expansion of open-vocabulary scene understanding to 3D scenes. Tschernezki et al. [25] introduced the DINO model [13] for extracting 2D features, which are then subsequently utilized for semantic feature learning within a trainable semantic field. Furthermore, Segment Anything Model (SAM) [6] has demonstrated remarkable zero-shot capabilities across a multitude of downstream 2D tasks. Extending this innovation, SAM3D [31] integrates the 2D masks generated by SAM with point clouds through projection, thereby imbuing the 3D point cloud space with semantic features.
@@ 3.1 Online Construction of Open-Vocabulary Field
Current methods about constructing dense open-set semantic scenes such as LERF [5] mainly rely on resource-intensive neural networks to embed CLIP features. This restricts their ability to be updated online. To build an online open-vocabulary mapping pipeline, we employ hierarchical scene representations of various scales and types to store geometric and semantic information about scene objects. Specifically, the scene is represented by three types of voxel grids: RGB, geometric, and semantic. These three distinct features are associated to each cell of the feature grid and updated in on online manner. We utilize Multilayer Perceptrons (MLPs) to decode occupancy, color and semantic information from these features and perform volume rendering. Specifically, for a spatial point $p$, we extract two fused latent vectors: the depth feature ${\phi}_{d}(p)$ and the color feature ${\phi}_{c}(p)$. We use trilinear interpolation between the eight neighboring voxels to extract latent vectors. These vectors are concatenated with the position-encoded point $p$ and fed into two separate MLPs, $f^{d}$ and $f^{c}$, to decode occupancy probability and color values, respectively.
@@ 4.1 Qualitative Results
O2V-mapping not only allows for the online construction of open-set vocabulary scenes but also significantly improves object-level scene understanding, semantic segmentation accuracy, and precision compared to LERF. For larger objects like "couch" that only appear in small portions in each image, O2V-mapping can still leverage the spatial consistency of implicit scene representations to understand the object as a whole, and demonstrate complete and clear object boundaries. For small yet complex objects like "potted plant", O2V-mapping can also describe precise object boundaries. Additionally, multiple objects of the same type, such as "chair" and "trash can" in the scene can also be assigned correct open-set semantic information by O2V-mapping. Even with slight differences in material, color, and shape among objects of the same type, they can still be clearly distinguished from surrounding objects of different types. For objects like "table" and "door" with uniform textures and flat surfaces in the scene, O2V-mapping exhibits excellent performance. We also tested our method on the real-world ScanNet dataset with depth noise, and the results in Fig. 5 demonstrate that our method has good robustness.
@@
Figure 5: Our method was tested on three scenes and 18 object types from the ScanNet dataset, and it still performed well in real scenes with depth noise.
@@ 4.2 Query Robustness
LERF is highly sensitive to query keywords, making it prone to false positives when the query keywords are ambiguous or exhibit a "bag-of-words" behavior and struggles to capture spatial relationships between objects. A common scenario is that when the queried object is not within the current viewpoint, LERF tends to exhibit a confused distribution of relevance scores (6). O2V-mapping exhibits robustness to query keywords. As shown in 6, in scenes where "door" is not present within the current viewpoint, there is no confused distribution of relevance scores.
@@ 5 Conclusion
Acknowledgements
This work is supported by Shanghai Pujiang Program (23PJ1400900) and State Key Laboratory of Intelligent Vehicle Safety Technology Open Fund Project (IVSTSKL-202317). Additionally, I extend my appreciation to the Shanghai Engineering Research Center of AI Robotics, Fudan University, China, and the Engineering Research Center of AI Robotics, Ministry of Education, China, for their invaluable support in the successful completion of this project.